OpenStudy (anonymous):

Solve. 4 – d < 4 + d A. d > –8 B. d > –4 C. d > 0 D. d > 8 Explain please but i think its d

OpenStudy (skullpatrol):

Why do you think it's d?

OpenStudy (skullpatrol):

$$\huge 4 – d < 4 + d$$

OpenStudy (anonymous):

because if you do all the following answers its wrong cuz its not less than right @skullpatrol im a little lost

OpenStudy (anonymous):

It is D because 4 - 8 = -4 and 4 + 8 = 12 -4 < 12

OpenStudy (anonymous):

yea thats exactly how i did it @Victoria_Davis_2018

OpenStudy (anonymous):

could you help with another one @Victoria_Davis_2018 The sum of two consecutive integers is at least 46. What is the least possible pair of integers? A. 21 and 23 B. 23 and 24 C. 22 and 23 D. 24 and 25 this one im completly lost

OpenStudy (skullpatrol):

What if I let d=1 Then 4-1=3 and 4 + 1 =5 3<5

OpenStudy (anonymous):

yea but thats not any of the following and y would wanna subtract 1 or add @skullpatrol

OpenStudy (anonymous):

22 + 24

OpenStudy (skullpatrol):

Your answer is d >8 I showed that d = 1 gives a true statement but is less than 8.

OpenStudy (anonymous):

ooooo i get it now thanxxx could help me with my other question i just put that one on this page to and that one i complety lost @skullpatrol

OpenStudy (skullpatrol):

Let's finish the first one by getting the correct answer.

OpenStudy (anonymous):

i dont get it @Victoria_Davis_2018

OpenStudy (anonymous):

I would say that the answer is 22 + 24 = 46 but that is not one of the choices

OpenStudy (anonymous):

i did get the correct answer and understand it :) @skullpatrol

OpenStudy (anonymous):

ik so what would i do im lost @Victoria_Davis_2018

OpenStudy (anonymous):

The answer is 23 + 24 = 47 because it has to equal AT LEAST 46 but he has to be the least possible pair to equal it or be greater

OpenStudy (anonymous):

@alvin78901 do you need anymore help?

OpenStudy (anonymous):

no im fine but thank you for your help @Victoria_Davis_2018